OpenStudy (anonymous):

The sum of five consecutive even integers is 0

OpenStudy (unklerhaukus):

-2,-1,0,1,2

OpenStudy (unklerhaukus):

ops evn only

myininaya (myininaya):

n n+1 n+2 n+3 n+4 there are 5 consecutive integers the sum means we have n+(n+1)+(n+2)+(n+3)+(n+4) we want this sum to be 0 so we have n+(n+1)+(n+2)+(n+3)+(n+4)=0 Collect like terms 5n+10=0 5n=-10 n=-2 if n=-2, then n+1=-1 n+2=0 n+3=1 n+4=2

myininaya (myininaya):

oops i missed that part too lol

OpenStudy (unklerhaukus):

-4,-2,0,2,4